OpenStudy (anonymous):

Find the product of the complex number and its conjugate. 1 + 2i

OpenStudy (anonymous):

complex cojugate is 1-2i and (1+2i)(1-2i) will be 1^2 - (2i)^2 = 1+4 = 5 since (i)^2 = -1

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@coolkid4 pleasure :)))))))))